susan is 5 years younger than her sister. the sum of their age is 51. define the variable
1
Expert's answer
2012-11-01T09:59:02-0400
Let age of Susan is x. Then age of sister is x+5. We can construct the following equation:
x+(x+5)=51
2x=51-5=46
x=23